# How General Are Measures of Choice Consistency? Evidence from Experimental and Scanner Data

### v1

@@ 3 Scanner Data and Experimental Design
In this section, we describe the scanner data and experiments. We examine purchase behaviors of consumers who shop at a leading retail company with more than 400 supermarkets spreading over 6 provinces in southern China. First, we obtained an individual transaction scanner dataset from the company to measure consistency scores in the field. A sample transaction dataset is presented in Table 1, where each observation represents the purchase of a single item, that is, one type of good bought by one consumer in a single shopping instance. The supermarket membership ID uniquely identifies each consumer. The dataset also includes the unique store ID, transaction timestamp, category of the good, the quantity purchased and the amount of expenditure. In addition, we conducted two lab-in-the-field experiments to measure consistency scores for risky, social, and food decisions. The combination of scanner and experimental datasets enables us to test the generalizability of measures of choice consistency.
@@
Membership ID Store ID Transaction timestamp Category Quantity (kg) Expenditure (RMB)
@@
3.1 Scanner Dataset
We focus on two categories of consumption, meat and vegetables, in the baseline analysis. The choice of these two categories is due to the following two reasons. First, because these two are most frequently purchased goods in the supermarkets, we are able to garner a sufficient number of observations. Second, because the unit for meat and vegetables is kilogram, it is convenient for us to have a refined measure of the quantity. We measure the quantity and price for meat and vegetables by averaging the quantity and prices of items in each category. We denote $q_{in}^{t}$ as the quantity of item $i$ in consumption category $I$ for consumer $n$ in month $t$, and $Q^{t}_{In}$ as the quantity of consumption category $I$ for consumer $n$ in month $t$, which is obtained by the sum of $q_{in}^{t}$. Q_In^t=∑_i∈I q_in^t. In a similar vein, we denote $P_{In}^{t}$ as the price of the consumption category $I$ for consumer $n$ in month $t$, which is obtained by averaging prices weighted by quantity: P_In^t=∑i∈IeintQInt, where $e_{in}^{t}$ is the expenditure for item $i$ in category $I$. Note that we use shelf price instead of the final price with discount (Echenique et al., 2011). With quantity $Q_{In}^{t}$ and price $P_{In}^{t}$, we can construct budget lines for consumer $n$ in month $t$ and perform revealed preference analysis for each consumer $n$ [^2] To perform the revealed preference analysis, we assume that preferences over meat and vegetables are weakly separable from other goods and services (Echenique et al., 2011; Dean and Martin, 2016).
It is worth noting that it is often challenging to construct individualized price using consumption data. For example, Echenique et al. (2011) assume the same price for each item for all consumers and Dean and Martin (2016) assume the same price for all items under the same good category for all consumers. Since we are able to access high frequency data at the individual level, we can construct individualized price index for each consumer. This helps to reduce measurement errors and increase the power of GARP tests.
We restrict our attention to 6,144 highly frequent consumers who have purchase records of the two categories for 24 consecutive months from January 2019 to December 2020. After matching the scanner dataset with participants in Experiment 1 introduced in the next subsection, we have a dataset of 1055 consumers, which we refer to as Scanner Dataset 1.
@@ 3.1 Scanner Dataset
3.2 Experiments
Figure 1: Screenshots for the Risk, Social and Food Tasks
Moreover, to help consumers in the supermarket understand the experimental tasks, we discretize the choice set as in Kim et al. (2018), by which they choose among 11 allocation options. [^3] Polisson and Quah (2013) show that Afriat’s theorem is applicable in discrete settings. Figure 1 presents a screenshot of the task. In total, participants face 22 decision tasks presented in random order. Different from Choi et al. (2007, 2014), we do not use randomly generated budget lines for each participant. Instead, every individual is presented with the same set of budget lines to ensure that the power for revealed preference tests is the same for all participants and to facilitate comparison across participants (Halevy et al., 2018). At the end of the experiment, we include a 10-question version of Big Five personality questions, seven questions from Raven’s progressive matrices, and collect demographic information in the post-experiment questionnaire.
To further investigate correlations of consistency scores across different tasks in the experiment, we conduct Experiment 2. In addition to risky decisions as in Experiment 1, we also measure consistency scores for social and food decisions. We use a modified dictator game to measure consistency scores for social decisions (Andreoni and Miller, 2002; Fisman et al., 2007). Participants allocate 100 tokens between themselves and another randomly matched participant across 22 rounds. In every round, each token has a different cash value for the individual and her matched participant.
To measure consistency scores for food decisions in the experiment, we construct a shopping environment (Chen et al., 2023; Harbaugh et al., 2001). In our design, participants allocate the expenditure of RMB 50 (USD 1 $\approx$ RMB 6.5) between a specific type of meat (ham) and a specific type of vegetables (tomato) in 22 rounds. The price level for each product is chosen from the range of actual prices from 2020 to 2021. We present the risk, social and food tasks in random order to participants and end the experiment with the post-experiment questionnaire identical to that in Experiment 1. Screenshots for the two experiments are presented in Figure 1, and detailed instructions are provided in Appendix C.
The two experiments were conducted from July to September 2021 and in December 2021, respectively. In Experiment 1, of the 6,144 frequent consumers identified in our baseline consumption analysis, we have contact information for around half of them. We invited these consumers to participate in the incentivized experiment in nearby supermarkets. In total, 1,055 highly frequent consumers participated in the experiment across 96 supermarkets. In Experiment 2, another 302 consumers participated in the experiment. [^4] To avoid learning spillover between experiments, we only invite consumers who did not participate in Experiment 1. Therefore, most of them are not highly frequent consumers and we do not have enough power to examine the consistency scores of their choices in the supermarket. All participants in both experiments received RMB 50 for completing the experiment. Additionally, we randomly paid 10 percent of participants based on their choices. Note that if the decision in food task was chosen for reward, we deposited redemption vouchers into the consumer’s membership account based on her choice. Each individual received an average payment of RMB 57.3 in Experiment 1 and RMB 55.1 in Experiment 2. Table A1 reports summary statistics for our participants in both experiments. For example, among the 1,055 participants in Experiment 1, the average monthly expenditure on meat and vegetables is RMB 482.2, with an average purchase frequency of 11.8 days per month.
@@ 4 Results
To examine the correlation of choice consistency between experimental and scanner data, we utilize choices made by 1,055 individuals both in Experiment 1 and Scanner Dataset 1. Figure 2 presents the cumulative distribution of consistency indices for risky decisions (solid line) and the summary statistics are in Table A2 (Panel A). Taking CCEI as the example, we find that 28.2% participants have no violations of GARP, and the average CCEI score is 0.941, which implies that the budget must be reduced by 5.9 percent to remove all GARP violations. [^5] The observed CCEI distribution is comparable to those in other experimental studies on risky decisions. For example, the average CCEI scores are 0.881 in Choi et al. (2014), 0.937 in Choi et al. (2007), and 0.979 in Halevy et al. (2018). To examine the power of the test, we calculate simulated scores by generating observations using uniformly random allocations over 11 options across 22 rounds for 1,000 times (Bronars, 1987). In contrast to the CCEI scores based on choice data, data from simulated random allocations (dashed line in Figure 2) have substantially lower scores ($p<0.01$, two-sided two-sample $t$-tests). [^6] Without specific clarification, we use two-sided tests throughout the study. For example, only 1.5 percent of their CCEI scores are over 0.9. This suggests that our choice data have sufficient power to detect GARP violations.
@@
Figure 3 presents the cumulative distribution of observed (solid line) and simulated (dashed line) scores of four consistency indices for consumption decisions in Scanner Dataset 1 (see Panel B in Table A2 for summary statistics). We find that 15.5% consumers have no violations of GARP, and the average CCEI score is 0.946.777Using scanner data, the average CCEI scores are 0.976 in Echenique et al. (2011), and 0.99 in Dean and Martin (2016). As supermarket consumers face different budget lines, we calculate simulated CCEI score for each consumer. Each simulated CCEI score is the average score generated by repeating the uniformly random allocations over budget shares along one consumer’s 24 budget lines for 100 times (Dean and Martin, 2016). The results show that simulated CCEI scores are significantly smaller than CCEI scores for actual choices ($p<0.01$, two-sample t-tests), and only 6.2 percent of them are over 0.9.888Using two-sample $t$-tests, all observed scores are significantly higher than simulated scores ($p<0.01$) in both Experiment 2 and Scanner Dataset 2. To avoid redundancy, we omit these repeated expressions in the following analyses. This confirms the sufficient power of budget lines in the scanner dataset.
@@
The consistency scores are uncorrelated between risky decisions in Experiment 1 and consumption decisions in Scanner Dataset 1.
@@
Given the above similar correlations among all consistency indices, we examine the relationships among these indices within each setting. Table 2 reports correlation coefficients for pairwise comparisons and shows that the average of correlation coefficients is 0.809. Whereas there have been theoretical discussions on how best to capture the extent of GARP violations (Apesteguia and Ballester, 2015; Echenique, 2021; Polisson and Quah, 2024), we show empirically that these indices are highly correlated. In subsequent analyses, for ease of presentation, we focus on results using CCEI and report results for other indices as robustness checks.
@@ 4.2 Experiment 2 & Scanner Dataset 2
In this section, we examine the correlations of choice consistency when making budgetary choices in Experiment 2 and consumption choices in Scanner Dataset 2. This analysis helps us see if the lack of correlation in consistency scores from risky decisions in Experiment 1 and consumption decisions in Scanner Dataset 1 is due to the different types of choices being made or different settings.
First, we calculate CCEI scores for each task in Experiment 2 (Figure A1 and Table A3) and their patterns are similar to those in Experiment 1. Next, we test pairwise correlations of CCEI scores across tasks (Figure 5). The results reveal that CCEI scores for risky decisions are moderately correlated with those both for social decisions (0.412, $p<0.01$) and food decisions (0.305, $p<0.01$) in Experiment 2, suggesting that consumers with higher consistency scores in the risk task tend to exhibit a higher level of consistency in the social and food tasks in the experiment. Meanwhile, the CCEI scores for social and food decisions are weakly correlated (0.254, $p<0.01$). These lead to our next observation.
@@
  • Notes: $N=822$ ($938$) for the three correlations across categories (time periods). Spearman’s rank correlations (r) are reported. The size of each circle represents the number of observations at the point. The lines represent the linear fit, and 95% confidence intervals are shown in the shaded area. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
We similarly summarize the consistency of choices in Scanner Dataset 2 (Figure A2 and Table A4), and examine correlations of consistency across non-overlapping pairs of consumption categories and time periods. [^9] We do not look at the correlations of consistency scores between pairs that share the same categories (time periods) because the same consumption choices in those overlapping categories (time periods) can create correlations. Figure 6 presents Spearman’s rank correlations of CCEI scores. These results indicate that the correlations are positive across most food categories, though the magnitude is low (0.173, $p<0.01$; -0.046. $p>0.1$; and 0.096, $p<0.01$). Similarly, there are also significantly positive but low correlations across different time windows (0.112, $p<0.01$; 0.186, $p<0.01$; and 0.143, $p<0.01$). We summarize these findings below.
Result 3.
The correlations of consistency scores are generally low both across consumption categories and over time periods in the supermarket.
@@ 4.3 Robustness Checks
In this subsection, we conduct several robustness checks on our main results including alternative indices of HMI, MPI, and MCI, the power of detecting GARP violations, imposing more assumptions on preferences, and fitness in structural estimations.
@@
Heterogeneous Power  Consumers face individualized budget sets in our scanner dataset, so the power for detecting GARP varies across consumers. To address this issue, we calculate Selton Score by subtracting simulated CCEI from the observed CCEI (Selten, 1991; Dean and Martin, 2016). In addition, we use the residual from a regression of simulated CCEI on observed CCEI.
@@
More (Realistic) Assumptions on Preferences  Since GARP measures the extent of maximizing some well-behaved, i.e., continuous and strictly increasing, utility function, it is possible that our choice dataset satisfies GARP, but may fail to meet requirements on particular preferences. Thus, we consider more realistic restrictions and compute new consistency indices under these conditions. Specifically, two properties related to income are often discussed. First, we examine the adherence to the normality of demand for both goods, which suggests that the demand for commodities increases as income rises, holding their prices constant (denoted as Normal) (Cherchye et al., 2018). [^10] Cherchye et al. (2018) provide the necessary and sufficient condition for the normality of both goods. To further assess how well our data align with this condition, we scale the budget in their NARP-I condition (Definition 3) by a factor $e$, and identify the largest $e$ that allows the dataset to meet the condition for both goods. We then regard this maximum $e$ as Normal. Second, we consider the homothetic preference, assuming a linear relationship between income level and consumption (Homothetic) (Heufer and Hjertstrand, 2019; Varian, 1983). Moreover, consumers’ utility may be negatively affected by the expenditure, especially in the supermarket with different expenditures across periods. Therefore, we consider the quasilinear utility $U(x)-px$, where the expenditure of goods reduces the utility in a linear form (Quasilinear) (Brown and Calsamiglia, 2007; Demetry et al., 2022). Meanwhile, to further evaluate choices on two contingent assets in the experiment, we also consider the compliance with first order stochastic dominance (FOSD) and expected utility (EU) maximization (Nishimura et al., 2017; Polisson et al., 2020).
@@
Preference Estimation  In addition to revealed preference analysis, we further measure the extent to which certain specific types of preferences fit the choice data. For risky decisions, we apply the disappointment aversion (DA) model, which is equivalent to rank dependent utility with binary states. For the other types of decisions, we use the constant elasticity of substitution (CES) utility functions (Fisman et al., 2007), in which parameter $\alpha$ captures the relative weight placed on the first good, and $\rho$ is the curvature of indifferent curves. We report the detailed estimation procedure in Appendix B. Because the log likelihood (LL) measures the goodness of fit of the utility model, we may interpret it as the degree of noise when making decisions.
@@
Exp. 1 vs. Scanner Dataset 1 Exp. 2 Scanner Dataset 2
@@
Selten Score -- -- --
Power-adjusted CCEI -- -- --
@@
LL
  • Notes: / / /  represents that the magnitude of correlation coefficient is in [0,0.1)/[0.1,0.3)/[0.3,0.5)/[0.5,1]. Specific values are detailed from Tables A5 to A7. -- denotes that the index is not applicable. Within Experiment 2, the three coefficients represent the correlations of consistency scores between risk and social tasks, risk and food tasks, and social and food tasks, respectively. In Scanner Dataset 2, the three coefficients across categories represent those between (Snacks,Meat) and (Veg.,Fruits), (Snacks,Veg.) and (Meat,Fruits), and (Snacks,Fruits) and (Meat,Veg.); those across time are correlations between (2018,2019) and (2020,2021), (2018,2020) and (2019,2021), and (2018,2021) and (2019,2020).
Table 3 summarizes the correlations using these consistency indices (see Table A5 to A7 for details). We find support for the robustness of Result 1 on no correlation between experimental and scanner data and Result 2 on moderate correlations within experimental data. Notably, after imposing more assumptions on consumption decisions in the supermarket, correlations within scanner data, as described in Result 3, increase substantially both in magnitude and significance.

### v2

@@ 3 Experimental Design and Scanner Data
In this section, we describe the experimental design and the scanner data. Our analysis focuses on consumers who shop at a leading retail company with more than 400 supermarkets spread over six provinces in southern China. We conducted two lab-in-the-field experiments using these consumers to measure consistency scores for risky, social, and food decisions. The combination of experimental and scanner datasets enables us to test the generalizability of measures of choice consistency.
3.1 Experiments
(a) Risk Task
(b) Social Task
(c) Food Task
Figure 1: Screenshots for the Risk, Social and Food Tasks
Moreover, to help consumers in the supermarket understand the experimental tasks, we discretize the choice set as in Kim et al. (2018), by which they choose among 11 allocation options. [^4] Polisson and Quah (2013) show that Afriat’s theorem is applicable in discrete settings. Figure 1(a) presents a screenshot of the task. In total, participants face 22 decision tasks presented in random order. Every individual is presented with the same set of budget lines to ensure that the power for revealed preference tests is the same for all participants and to facilitate comparison between participants (Halevy et al., 2018). At the end of the experiment, we include a 10-question version of the Big Five personality questions, seven questions from Raven’s progressive matrices, and collect demographic information in the post-experiment questionnaire.
To further investigate the correlations of consistency scores between different tasks in the experiment, we conduct Experiment 2. In addition to risky decisions as in Experiment 1, we also measure consistency scores for social and food decisions. We use a modified dictator game to measure consistency scores for social decisions (Andreoni and Miller, 2002; Fisman et al., 2007). Participants allocate 100 tokens between themselves and another randomly matched participant in 22 rounds. In each round, each token has a different cash value for the individual and her matched participant.
To measure consistency scores for food decisions in the experiment, we construct a shopping environment (Chen et al., 2023; Harbaugh et al., 2001). In our design, participants allocate the expenditure of RMB 50 (USD 1 $\approx$ RMB 6.5) between a specific type of meat (ham) and a specific type of vegetables (tomato) in 22 rounds. The price level for each product is chosen from the range of actual prices from 2020 to 2021. We present the risk, social, and food tasks in random order to participants and end the experiment with the post-experiment questionnaire identical to that in Experiment 1. Screenshots for the two tasks are presented in Figures 1(b) and 1(c), and detailed instructions are provided in Appendix C.
The two experiments were conducted from July to September 2021 and in December 2021, respectively. In Experiment 1, of the 6,144 consumers identified as highly frequent consumers introduced in the next subsection, contact information is available for approximately half of them. We invited these consumers to participate in the incentivized experiment in nearby supermarkets. In total, 1,055 highly frequent consumers participated in the experiment in 96 supermarkets. In Experiment 2, another 302 consumers participated in the experiment. To avoid learning spillover between experiments, we only invited consumers who did not participate in Experiment 1. Therefore, most of them are not highly frequent consumers and we do not have enough power to examine the consistency scores of their choices in the supermarket.
All participants in both experiments received RMB 50 for completing the experiment. Additionally, we randomly paid 10 percent of participants based on their choices. Note that if the decision in the food task was chosen as reward, we deposited redemption vouchers into the consumer’s membership account based on her choices. Each individual received an average payment of RMB 52.5 in Experiment 1 and RMB 52.0 in Experiment 2. In Experiment 2, participants also received vouchers redeemable on average for 0.04 kg of hams and 0.10 kg of tomatoes. Table A1 reports the summary statistics for our participants in both experiments. For example, among the 1,055 participants in Experiment 1, the average monthly expenditure on meat and vegetables is RMB 482.2, with an average purchase frequency of 11.8 days per month.
3.2 Scanner Dataset
We obtained a transaction-level scanner dataset from the company to measure consistency scores in the field. A sample transaction dataset is presented in Table 1, where each observation represents the purchase of a single item, that is, one type of good bought by a consumer in a single shopping instance. The supermarket membership ID uniquely identifies each consumer. The dataset also includes the unique store ID, transaction timestamp, category of the goods, the quantity purchased, the amount of expenditure, and additional transaction details such as discounts.
@@ 3.2 Scanner Dataset
Membership
ID
Store
ID
Transaction
timestamp
Category
Quantity
Expenditure
@@
We now introduce how we construct individual-level choice dataset $\{(p^{t},x^{t})\}_{t=1}^{T}$ based on supermarket purchases. We focus on two categories of consumption, meat and vegetables, in the baseline analysis. The choice of these two categories is due to the following two reasons. First, because these two are most frequently purchased goods in the supermarkets, we are able to garner a sufficient number of observations. Second, because the unit for meat and vegetables is kilogram, it is convenient for us to have a refined measure of the quantity. We measure the quantity and price for meat and vegetables by averaging the quantity and prices of items in each category. We denote $q_{in}^{t}$ as the quantity of item $i$ in consumption category $I$ for consumer $n$ in month $t$, and $Q^{t}_{In}$ as the quantity of consumption category $I$ for consumer $n$ in month $t$, which is obtained by the sum of $q_{in}^{t}$.
$Q_{In}^{t}=\sum_{i\in I}q_{in}^{t}.$
In a similar vein, we denote $P_{In}^{t}$ as the price of the consumption category $I$ for consumer $n$ in month $t$, which is obtained by averaging prices weighted by quantity:
$P_{In}^{t}=\frac{\sum_{i\in I}e_{in}^{t}}{Q_{In}^{t}},$
where $e_{in}^{t}$ is the expenditure for item $i$ in category $I$. With quantity $Q_{In}^{t}$ and price $P_{In}^{t}$, we can construct budget lines for consumer $n$ in month $t$ and perform revealed preference analysis for each consumer $n$.
The construction of budget lines involves certain assumptions and simplifications. First, our approach of forming composite goods and aggregating purchases over time aligns with standard practices in applied demand analysis (e.g., Barten, 1969; Deaton and Muellbauer, 1980; Allen et al., 2024). Second, utilizing high-frequency individual-level data, we focus on consumers with consecutive purchase records and create an individualized price index for each consumer. This contrasts with studies like Echenique et al. (2011), which assume the same price for each item for all consumers, and Dean and Martin (2016), which assume uniform pricing for all items within a category. Many other studies adopt similar “same-price” assumptions (e.g., Allen and Rehbeck, 2020; Demuynck and Seel, 2018; Tipoe, 2021). Our individualized price index strengthens the power of GARP tests and reduces measurement errors. Third, we assume preferences for meat and vegetables are weakly separable from other goods and services, a common assumption in applied demand analysis (e.g., Blundell et al., 2003; Dean and Martin, 2016; Deaton and Muellbauer, 1980; Echenique et al., 2011). This enables independent examination of consumer choices within these categories. In Section 4.3, we explore the impact of other goods and income using some revealed preference techniques (Brown and Calsamiglia, 2007; Deb et al., 2023; Cherchye et al., 2018). Lastly, we use the final price with discounts to reduce measurement errors, and use shelf prices as a robustness check, with our main results remaining consistent.
Following the above construction of budget lines, we identify 6,144 high-frequency consumers (as noted in Section 3.1) who have purchase records in both categories for 24 consecutive months from January 2019 to December 2020. After matching the scanner data with participants from Experiment 1, we obtain a final dataset of 1,055 consumers, which we refer to as Scanner Dataset 1.
@@ 4 Results
To examine the correlation of choice consistency between experimental and scanner data, we utilize choices made by 1,055 individuals both in Experiment 1 and Scanner Dataset 1. Figure 2 presents the cumulative distribution of consistency indices for risky decisions (solid line) and the summary statistics are in Table A2 (Panel A). Taking CCEI as an example, we find that 28.2% participants do not have GARP violations, and the average CCEI score is 0.941, which implies that the budget must be reduced by 5.9 percent to remove all GARP violations. The observed CCEI distribution is comparable to those in other experimental studies on risky decisions. For example, the average CCEI scores are 0.881 in Choi et al. (2014), 0.937 in Choi et al. (2007), and 0.979 in Halevy et al. (2018). To examine the power of the test, we calculate simulated scores by generating observations using uniformly random allocations over 11 options in 22 rounds for 1,000 times (Bronars, 1987). In contrast to the CCEI scores based on choice data, the data from the simulated random allocations (dashed line in Figure 2) have substantially lower scores ($p<0.01$, two-sided two-sample $t$-tests). For example, only 1.5 percent of their CCEI scores are greater than 0.9. This suggests that our choice data have sufficient power to detect GARP violations.
@@
Figure 3 presents the cumulative distribution of observed (solid line) and simulated (dashed line) scores of four consistency indices for consumption decisions in Scanner Dataset 1 (see Panel B in Table A2 for summary statistics). We find that 15.5% consumers have no GARP violations, and the average CCEI score is 0.946. Using scanner data, the average CCEI scores are 0.976 in Echenique et al. (2011), and 0.99 in Dean and Martin (2016). As supermarket consumers face different budget lines, we calculate the simulated CCEI score for each consumer. Each simulated CCEI score is the average score generated by repeating uniformly random allocations over budget shares along the 24 budget lines of a consumer 100 times (Dean and Martin, 2016). The results show that simulated CCEI scores are significantly smaller than CCEI scores for actual choices ($p<0.01$, two-sample t-tests), and only 6.2 percent of them are over 0.9.555Using two-sample $t$-tests, all observed scores are significantly higher than simulated scores ($p<0.01$) in both Experiment 2 and Scanner Dataset 2. To avoid redundancy, we omit these repeated expressions in the following analyses. This confirms the sufficient power of budget lines in the scanner dataset.
@@
The consistency scores are not correlated between risky decisions in Experiment 1 and consumption decisions in Scanner Dataset 1.
@@
Given the above similar correlations among all consistency indices, we examine the relationships among these indices within each setting. Table 2 reports the correlation coefficients for pairwise comparisons and shows that the average of the correlation coefficients is 0.809. Although there have been theoretical discussions on how best to capture the extent of GARP violations (Apesteguia and Ballester, 2015; Echenique, 2021; Polisson and Quah, 2024), we show empirically that these indices are highly correlated. In subsequent analyses, for ease of presentation, we focus on results using CCEI and report results for other indices as robustness checks.
@@ 4.2 Experiment 2 & Scanner Dataset 2
In this section, we examine the correlations of choice consistency when making budgetary choices in Experiment 2 and consumption choices in Scanner Dataset 2. This analysis helps us see whether the lack of correlation in consistency scores between risky decisions in Experiment 1 and consumption decisions in Scanner Dataset 1 is due to the different types of choices being made or different settings.
First, we calculate the CCEI scores for each task in Experiment 2 (Figure A1 and Table A3) and their patterns are similar to those in Experiment 1. Next, we test pairwise correlations of the CCEI scores between tasks (Figure 5). The results reveal that the CCEI scores for risky decisions are moderately correlated with those for both social decisions (0.412, $p<0.01$) and food decisions (0.305, $p<0.01$) in Experiment 2, suggesting that consumers with higher consistency scores in the risk task tend to exhibit a higher level of consistency in the social and food tasks in the experiment. Meanwhile, the CCEI scores for the social and food decisions are weakly correlated (0.254, $p<0.01$). These lead to our next observation.
@@
We similarly summarize the consistency of choices in Scanner Dataset 2 (Figure A2 and Table A4), and examine correlations of consistency across nonoverlapping pairs of consumption categories and time periods. Note that we do not look at the correlations of consistency scores between pairs that share the same categories (time periods) because the same consumption choices in those overlapping categories (time periods) can create correlations. Figure 6 presents Spearman’s rank correlations of CCEI scores. The correlations across categories are generally positive and statistically significant (0.173, $p<0.01$; -0.046. $p>0.1$; and 0.096, $p<0.01$). Similarly, the correlations across time windows are consistently positive and significant (0.112, $p<0.01$; 0.186, $p<0.01$; and 0.143, $p<0.01$). We summarize these findings in the following.
Result 3.
The consistency scores show generally positive and significant correlations across consumption categories and time periods in the supermarket, although the magnitude of these correlations is relatively low.
@@
  • Notes: $N=822$ (938938) for the three correlations across categories (time periods). Spearman’s rank correlations (r) are reported. The size of each circle represents the number of observations at the point. The lines represent the linear fit, and 95% confidence intervals are shown in the shaded area. * $p<0.10$, ** $p<0.05$, *** $p<0.01$.
@@ 4.3 Robustness Checks
In this subsection, we conduct several robustness checks on our main results, including alternative indices of HMI, MPI, and MCI, the power to detect GARP violations, imposing more assumptions on preferences, and fitness in structural estimations.
@@
Heterogeneous Power Consumers face individualized budget sets in our scanner dataset, so the power to detect GARP varies across consumers. To address this issue, we calculate the Selten score by subtracting the simulated CCEI from the observed CCEI (Selten, 1991; Dean and Martin, 2016). In addition, we use the residual from a regression of simulated CCEI on observed CCEI (Power-adjusted CCEI).
@@
More (Realistic) Assumptions on Preferences Since GARP measures the extent of maximizing some well-behaved, i.e., continuous and strictly increasing, utility function, it is possible that our choice dataset satisfies GARP but may fail to meet requirements on particular preferences. Thus, we consider more realistic restrictions and compute new consistency indices under these conditions. Specifically, two properties related to income are often discussed. First, we examine the adherence to the normality of demand for both goods, suggesting that the demand for commodities increases as income increases, keeping their prices constant (denoted as Normal) (Cherchye et al., 2018). [^6] Cherchye et al. (2018) provide the necessary and sufficient condition for the normality of both goods. To further assess how well our data align with this condition, we scale the budget in their NARP-I condition (Definition 3) by a factor $e$, and identify the largest $e$ that allows the dataset to meet the condition for both goods. We then regard this maximum $e$ as Normal. Second, we consider the homothetic preference, assuming a linear relationship between income level and consumption (Homothetic) (Heufer and Hjertstrand, 2019; Varian, 1983). Moreover, consumers’ utility may be negatively affected by the expenditure, especially in the supermarket with different expenditures across periods. Therefore, we consider the quasilinear utility U(x)pxU(x)-px, where the expenditure of goods reduces the utility in a linear form (Quasilinear) (Brown and Calsamiglia, 2007; Demetry et al., 2022). Meanwhile, to further evaluate choices on two contingent assets in the experiment, we also consider the compliance with first-order stochastic dominance (FOSD) and expected utility (EU) maximization (Nishimura et al., 2017; Polisson et al., 2020).
@@
Preference Estimation In addition to the revealed preference analysis, we further measure the extent to which certain specific types of preferences fit the choice data. For risky decisions, we apply the disappointment aversion (DA) model, which is equivalent to rank-dependent utility with binary states. For the other types of decision, we use the utility functions of constant elasticity of substitution (CES) (Fisman et al., 2007), in which the parameter $\alpha$ captures the relative weight placed on the first good, and $\rho$ is the curvature of indifferent curves. We report the detailed estimation procedure in Appendix B. Because the log likelihood (LL) measures the goodness of fit of the utility model, we may interpret it as the degree of noise when making decisions.
Table 3 summarizes the correlations using these consistency indices (see Tables A5 to A7 for details). We find support for the robustness of Result 1 on no correlation between experimental and scanner data and Result 2 on moderate correlations within experimental data. Notably, after imposing more assumptions on preferences or evaluating the fit of preference structural estimation in the supermarket, correlations within scanner data, as described in Result 3, increase substantially both in magnitude and significance. Therefore, despite the inherent challenges in constructing consumption choices and budget constraints from scanner data, we can recover certain aspects of choice consistency in supermarkets.
@@
Exp. 1 vs. Scanner
Dataset 1
Exp. 2 Scanner Dataset 2
@@
Selten score - - -
Power-adjusted CCEI - - -
@@
$L$
  • Notes: / / / represents that the magnitude of correlation coefficient is in [0,0.1)/[0.1,0.3)/[0.3,0.5)/[0.5,1]. Specific values are detailed from Tables A5 to A7. - denotes that the index is not applicable. Within Experiment 2, the three coefficients represent the correlations of consistency scores between risk and social tasks, risk and food tasks, and social and food tasks, respectively. In Scanner Dataset 2, the three coefficients across categories represent those between (Snacks,Meat) and (Veg.,Fruits), (Snacks,Veg.) and (Meat,Fruits), and (Snacks,Fruits) and (Meat,Veg.); those across time are correlations between (2018,2019) and (2020,2021), (2018,2020) and (2019,2021), and (2018,2021) and (2019,2020).
Additionally, we also conduct a few more robustness checks. First, we use shelf prices instead of final price with discounts. Second, we restrict our analysis to the most frequently purchased subcategory, i.e., pork and leaf vegetables, within each category and calculate the consistency scores accordingly. Third, we assume that the price in each category is the same for all consumers in each month (Echenique et al., 2011; Dean and Martin, 2016). Specifically, we obtain the aggregate price index $P_{I}^{t}$ for category $I$ in month $t$: $P_{I}^{t}=\frac{\sum_{i\in I}e_{i}^{t}}{Q_{I}^{t}}$, where $e_{i}^{t}$ represents the total expenditure of all consumers on the item $i$ in category $I$ in month $t$, and $Q_{I}^{t}$ is the total quantity of items purchased in category $I$ in month $t$. In these analyses, all correlation coefficients are close to our baseline findings.